A bookcase is to have 4 shelves including the top as pictured below.

The width is to be 7 feet less than 2 times the height. Find the width and the height if the carpenter expects to use 22 feet of lumber to make it.
Width: feet
Height: feet

Answer by macston(5194)

There is no picture below, so this may not be correct:
H=height; W=width=2H-7ft
W=2H-7ft
2H+4W=22ft Substitute for W
2H+4(2H-7ft)=22ft
2H+8H-28ft=22ft Add 28 ft to each side.
10H=50ft
H=5 ft ANSWER 1: The bookshelf is 5 feet high.
W=2H-7ft=2(5ft)-7ft=10ft-7ft=3ft ANSWER 2: the bookshelf is 3 feet wide.
CHECK:
2H+4W=22ft
2(5ft)+4(3ft)=22ft
10ft+12ft=22ft
22ft=22ft